# Data-Efficient 3D Visual Grounding via Order-Aware Referring

### v3

@@ Abstract
3D visual grounding aims to identify the target object within a 3D point cloud scene referred to by a natural language description. Previous works usually require significant data relating to point color and their descriptions to exploit the corresponding complicated verbo-visual relations. In our work, we introduce Vigor, a novel Data-Efficient 3D Visual Grounding framework via Order-aware Referring. Vigor leverages LLM to produce a desirable referential order from the input description for 3D visual grounding. With the proposed stacked object-referring blocks, the predicted anchor objects in the above order allow one to locate the target object progressively without supervision on the identities of anchor objects or exact relations between anchor/target objects. In addition, we present an order-aware warm-up training strategy, which augments referential orders for pre-training the visual grounding framework. This allows us to better capture the complex verbo-visual relations and benefit the desirable data-efficient learning scheme. Experimental results on the NR3D and ScanRefer datasets demonstrate our superiority in low-resource scenarios. In particular, Vigor surpasses current state-of-the-art frameworks by 9.3% and 7.6% grounding accuracy under 1% data and 10% data settings on the NR3D dataset, respectively.
@@ 1 Introduction
To eliminate the need of human annotation for training grounding models, recent approaches [18, 5] leverage large language models (LLMs) for automatic dissection of the descriptions and generation of prior linguistic knowledge. For example, NS3D [18] utilizes Codex [10] to parse descriptions into nested expressions and designs a neuro-symbolic framework to find the target object step-by-step. However, it only considers fixed-template relations between objects (e.g., below/above, near/far, etc.) and cannot be easily extended to arbitrary natural descriptions. Inspired by the mechanism of human perception system [32, 9], CoT3DRef [5] generates the referential order of a description that points from anchor objects to the final target object using LLM. For example, for a description “Find the water bottle on the table nearest to the door.”, the referential order is generated as { “door” (anchor), “table” (anchor), “water bottle” (target) }. Additionally, it utilizes a rule-based algorithm to localize the identities of the above anchor/target objects, which guides a transformer-based module to predict the final target object. However, as noted in [5], such rule-based identity prediction might not be applicable for scenarios with complex language descriptions.
@@ 3 Methodology
3.1 Problem Formulation and Framework Overview
@@ 4 Experiments
4.1 Dataset
NR3D
NR3D [2] dataset consists of 707 indoor scenes in ScanNet [12] with 28715/7485 description-target pairs in the training/testing set, where the descriptions are collected from human annotators. There are 524 different object classes in the scenes in total. NR3D assumes perfect class-agnostic object proposals, where each point in the scene is properly assigned to its corresponding proposal. As a result, models are only required to classify the target object that uniquely matches the description among all proposals in the scene, with classification accuracy (Acc in %) being the metric. Please refer to the supplementary material for the details of the implementation (A) and experiments on an additional dataset ScanRefer [8] B.2.1.
4.2 Quantitative Results for Data Efficiency
@@
We present the quantitative results on NR3D, with consideration of different amounts of available training samples against several baselines by reproducing from their official implementation. As shown in Table 1, Vigor possesses a considerably superior performance when training with 1% (287), 2.5% (717), 5% (1435), and 10% (2871) NR3D training set samples. Specifically, using only 1% data, Vigor achieves 33.5 overall Acc, which even surpasses SOTA methods with 10% of data. This suggests that Vigor is preferable for 3D grounding, especially when paired training data is limited.
Method Hard Easy View-Dep. View-Indep. Overall
Referit3D [2] 19.5 27.3 21.2 24.2 23.3
TransRefer3D [17] 21.6 29.9 22.9 27.0 25.7
SAT [46] 22.5 27.6 21.7 26.6 25.0
BUTD-DETR [21] 25.9 41.9 29.1 34.8 33.3
MVT [20] 22.9 30.3 25.4 27.1 26.5
MVT + CoT3DRef [5] 32.7 43.2 34.0 39.8 37.9
ViL3DRel + CoT3DRef [5] 32.4 44.7 33.4 40.9 38.4
Vigor (Ours) 39.1 53.3 45.3 46.4 46.0
4.3 Ablation Studies
We ablate different components of our Vigor in Table 3, with performances under 1%, 10%, and 100% training samples available to explore the effectiveness of each component with different amounts of data. Baseline model A extracts class names of anchor/target objects from $D$ using a language parser [6] and constructs the referential order according to the appearance of the names in $D$ directly. Also, we employ each of our Referring Blocks $R_{i}$ in model A without the FE module, i.e., directly using $F_{i}$ to calculate attention with text features of $O_{i:B}$ and $D$. Model B enhances the accuracy on 1% and 10% training data by conducting our two-stage object ordering with LLM. When further applying the order-aware pre-training in model C, significant improvements in all settings, especially for 1% available data, are observed. The scarcity of training data often impedes the ability of the model to learn the concept of order and relations among objects implicitly. By conducting our pre-training strategy, our Vigor is able to learn foundational concepts of ordering and relations to locate the target objects progressively. Finally, our full model in the last row, incorporating FE modules each $R_{i}$ to enhance features of anchor/target objects with $F^{mask}_{i}$, achieves optimal results on both settings. This verifies the success of our proposed modules and warm-up strategy, especially when the available training data is very limited.
@@ A.1 Pre-Training Sample Synthesis and Training Pipeline
This section provides the pseudocode of synthesizing order-aware samples for Vigor’s pre-training in Sec. 3.4 and the complete training pipeline in Sec. 3.5 in our main paper, respectively. Specifically, Algorithm A1 demonstrates the pipeline to synthesize an order-aware pre-training sample given object proposals $P$ and predicted object labels $L$. On the other hand, A2 illustrates the complete pipeline to train Vigor with synthesized samples and natural-description samples, such as NR3D and ScanRefer.
Algorithm A1 Order-Aware Sample synthesis for Vigor Pre-training
@@ A.2 Implementation and Details and Experimental Settings
Algorithm A2 Vigor Training Pipeline
@@
6:    use A1 with $\{P,L,B\}$ as inputs to synthesize $\{D^{aug},O^{aug}_{1:B},p^{aug}_{1:B}\}$.
@@ A.2.1 NR3D
A.2.2 Baselines
For baselines [2, 17, 46, 21, 20, 5, 8, 53, 31, 52, 44] on NR3D and ScanRefer in Table 1,  2, and  3, we utilize their official public implementations with different amounts of available training samples and the full testing set to evaluate their low-resource performance. For the full-data ($100$%) scenario, we acquire their performance either on the official leaderboard of NR3D/ScanRefer or their published papers. Especially, the implementation of CORE-3DVG [44] is currently unavailable, so we only report its full-data ScanRefer performance, which is presented in their paper.
@@ B.1 More Results on NR3D
Figure A5: Quantitative results on NR3D We can see that when the amount of data is relatively many (above 30%), our Vigor is comparable to MVT+CoT3DRef [5]. However, as the amount of data reduces, our Vigor performs better over MVT-CoT3DRef.
In Sec. 4, we have shown comparisons between Vigor and several state-of-the-arts under data efficient scenarios (1% $\sim$ 10% of data). Here, we provide a comparison between Vigor and MVT+CoT3DRef [5] from 1% to 100% of data in Figure A5. In this figure, Vigor outperforms CoT3DRef with a noticeable margin when the amount of data is limited and is comparable to CoT3DRef when the data amount is over 30%, showing that our Vigor is suitable under different settings.
B.2 Experiments on ScanRefer
B.2.1 Dataset and setting
ScanRefer [8] contains 36665/9508 description-target pairs across a total of 800 indoor scenes in its training/validation set, where the descriptions are also collected from human annotators. Also derived from ScanNet [12] but different from NR3D, perfect object proposals are not available in ScanRefer, and therefore, additional object proposers are required for all methods. Nevertheless, Acc in % under 0.25 and 0.5 intersection over union (IoU) are used as the metrics for ScanRefer. In ScanRefer, since ground-truth object proposals are unavailable, we adopt the visual encoder of M3DRef-CLIP [52] that adopts PointGroup [22] to perform object segmentation as proposals $P$ and object classification as labels $L$. However, we do not perform object-aware pre-training for ScanRefer since imperfect object proposals may lead to noisy synthetic samples and hinder Vigor’s training stability.
B.2.2 Quantitative Results
The results are shown in Table A4. We can see that our Vigor demonstrates slightly preferable results with only 5% and 10% of data compared to other approaches. It is comparable to SOTA methods when given 100% data. Note that we only report CORE-3DVG’s [44] results on 100% (36665) ScanRefer training samples since its implementation is currently unavailable.
Table A4: Grounding accuracy (%) on the ScanRefer validation set. In this table, different amounts of training data are considered. Note that the official implementation of CORE-3DVG [44] has not yet been released; we show the numbers in their original paper in this table.
@@ B.2.2 Quantitative Results
5% 10% 100%
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
ScanRefer [8] 23.0 12.0 27.4 15.0 35.5 22.4
3DVG-Trans. [53] 35.3 23.3 39.0 29.0 45.9 34.5
3D-SPS [31] 28.4 16.9 32.9 22.9 48.8 37.0
BUTD-DETR [21] 38.2 26.2 40.3 28.3 52.2 39.8
M3DRef-CLIP [52] 37.2 29.9 40.0 32.4 52.5 44.7
CORE-3DVG [44] - - - - 56.8 43.8
Vigor (Ours) 39.8 31.5 43.6 34.9 52.8 44.8
@@ Appendix C Ablation Studies on Training Objectives
This section performs the ablation study on several training objectives used for Vigor mentioned in Sec. 3. In particular, we investigate the influence of $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ for low-resource (1% training data) and full-resource (100% training data) scenarios, as shown in A5. Note that we only conduct target object classification on the text feature $T$ when deactivating $\mathcal{L}_{text}$, and apply classification on both anchor and target objects when adopting $\mathcal{L}_{text}$, as mentioned in Sec. 3.3. Also, for all experiments in A5, the pre-training with synthetic data in Sec. 3.4 is applied for fair comparison. In summary, $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ all impose positive effects on models’ performances under the two settings, verifying Vigor’s design in Sec. 3.
Table A5: Ablation studies on training objectives. Note that for all ablated settings, the proposed order-aware pre-training in Sec. 3.4 is applied for fair comparison.
@@ Appendix D Model Performance Under Different Order Length
Figure A6: (a) Grounding accuracies with varying Object-Referring block numbers $B$ and (b) statistics of referential order length of NR3D. It can be seen that the model performance saturates at $B$=4, matching length statistics of NR3D.
In this section, we analyze the effectiveness of padded order length $B$ of $O_{1:B}$ (and also the number of Referring Blocks) in Fig. A6, which illustrates the performances with different $B$ using 1% and 100% training data in Fig. 6(a). We also show the statistics of the original order length generated by our two-stage referential ordering in Fig. 6(b). We observe that although performances gradually increase with larger $B$ (i.e., considering more potential anchor objects appeared in $D$ achieves better accuracy), the performance gain saturates at $B=4$. This can be explained by looking at Fig. 6(b), which shows that only a very small amount of data exceeds an order length of $4$ in both training and testing pairs. This suggests that our selection of $B=4$ is reasonable, optimally balancing computational efficiency with prediction accuracy.
Table A6: Grounding accuracy on NR3D subsets regarding different parsed referential order lengths. Note that referential order length $=1$ means only the target object is mentioned in the description. The improvement of Vigor over MVT [20] grows as the parsed order length increases.
@@
testing set) have 1 or 2 anchor objects mentioned other than the target object. Similarly, samples with an order length of 4 or 5 (209 samples in the testing set) have 3 or 4 anchor objects mentioned other than the target object. A longer referential order may generally denote a longer and more complicated description. As shown in A6, though only a 2.2% performance gain is obtained for target-only samples, Vigor is more advantageous when dealing with lengthy descriptions, with 5.3% performance gain achieved for descriptions with order length of 4 or 5.

### v4

@@ Abstract
3D visual grounding aims to identify the target object within a 3D point cloud scene referred to by a natural language description. Previous works usually require significant data relating to point color and their descriptions to exploit the corresponding complicated verbo-visual relations. In our work, we introduce Vigor, a novel Data-Efficient 3D Visual Grounding framework via Order-aware Referring. Vigor leverages LLM to produce a desirable referential order from the input description for 3D visual grounding. With the proposed stacked object-referring blocks, the predicted anchor objects in the above order allow one to locate the target object progressively without supervision on the identities of anchor objects or exact relations between anchor/target objects. We also present an order-aware warm-up training strategy, which augments referential orders for pre-training the visual grounding framework, allowing us to better capture the complex verbo-visual relations and benefit the desirable data-efficient learning scheme. Experimental results on the NR3D and ScanRefer datasets demonstrate our superiority in low-resource scenarios. In particular, Vigor surpasses current state-of-the-art frameworks by 9.3% and 7.6% grounding accuracy under 1% data and 10% data settings on the NR3D dataset, respectively.
@@ 1 Introduction
To eliminate the need of human annotation for training grounding models, recent approaches [59, 43, 18, 5] leverage pre-trained 2D priors (e.g., SAM [25], LDM [38]) or large language models (LLMs) for automatic dissection of the descriptions and generation of prior linguistic knowledge. For example, Diff2Scene [59] conducts the use of LDM to obtain text-conditioned 2D semantic maps as pseudo labels to guide a 3D segmentation model to achieve scene understanding and produce zero-shot 3D visual grounding. Unfortunately, limited by the spatial understanding ability of 2D diffusion models between multiple objects as described in [33], the ability of Diffi2Scene to achieve effective visual grounding for complex description with multiple anchor objects is still unclear. On the other hand, NS3D [18] utilizes Codex [10] to parse descriptions into nested expressions and designs a neuro-symbolic framework to find the target object step-by-step. However, it only considers fixed-template relations between objects (e.g., below/above, near/far, etc.) and cannot be easily extended to arbitrary natural descriptions. Inspired by the mechanism of human perception system [34, 9], CoT3DRef [5] generates the referential order of a description that points from anchor objects to the final target object using LLM. For example, for a description “Find the water bottle on the table nearest to the door.”, the referential order is generated as { “door” (anchor), “table” (anchor), “water bottle” (target) }. Additionally, it utilizes a rule-based algorithm to localize the identities of the above anchor/target objects, which guides a transformer-based module to predict the final target object. However, as noted in [5], such rule-based identity prediction might not be applicable for scenarios with complex language descriptions.
@@ 3 Methodology
3.1 Problem Formulation and Model Overview
@@ 4 Experiments
4.1 Dataset
NR3D
NR3D [2] dataset consists of 707 indoor scenes in ScanNet [12] with 28715/7485 description-target pairs in the training/testing set, where the descriptions are collected from human annotators. There are 524 different object classes in the scenes in total. NR3D provides ground-truth class-agnostic object proposals, where each point in the scene is properly assigned to its corresponding proposal. As a result, models are only required to classify the target object that uniquely matches the description among all proposals in the scene, with classification accuracy (Acc in %) being the metric.
ScanRefer
ScanRefer [8] contains 36665/9508 description-target pairs across a total of 800 indoor scenes in its training/validation set, where the descriptions are also collected from human annotators. Also derived from ScanNet [12] but different from NR3D, perfect object proposals are not available in ScanRefer, and therefore, additional object proposers are required for all methods. Nevertheless, Acc in % under 0.25 and 0.5 intersection over union (IoU) are used as the metrics for ScanRefer. In ScanRefer, since ground-truth object proposals are unavailable, we adopt the visual encoder of M3DRef-CLIP [56] that applies PointGroup [22] to perform object segmentation as proposals $P$ and object classification as labels $L$. We do not perform object-aware pre-training for ScanRefer since imperfect object proposals may lead to noisy synthetic samples and hinder Vigor’s training stability.
Table 2: Grounding accuracy (%) on the ScanRefer validation set. In this table, different amounts of training data are considered.
Method Labeled Training Data
5% 10%
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
ScanRefer [8] 23.0 12.0 27.4 15.0
3DVG-Trans. [57] 35.3 23.3 39.0 29.0
3D-SPS [32] 28.4 16.9 32.9 22.9
BUTD-DETR [21] 38.2 26.2 40.3 28.3
M3DRef-CLIP [56] 37.2 29.9 40.0 32.4
Vigor (Ours) 39.8 31.5 43.6 34.9
4.2 Quantitative Results for Data Efficiency
We present the quantitative results on NR3D and ScanRefer, with consideration of different amounts of available training samples against several baselines by reproducing from their official implementation. Results of NR3D are shown in Table 1. Vigor possesses a considerably superior performance when training with 1% (287), 2.5% (717), 5% (1435), and 10% (2871) NR3D training set samples. Specifically, using only 1% data, Vigor achieves 33.5 overall Acc, which even surpasses SOTA methods with 10% of data. This suggests that Vigor is preferable for 3D grounding, especially when paired training data is limited. Table 2 shows the results on ScanRefer datasets with 5% (1833) and 10% (3666) training samples. Vigor still surpasses all baselines under the conditions where imperfect object proposals are used, indicating the robustness of our methods. The detailed performance on different official subsets of NR3D are in the Supp. B.1.
@@ 4.2 Quantitative Results for Data Efficiency
4.3 Ablation Studies
We ablate different components of our Vigor in Table 3, with performances on NR3D under 1%, 10%, and 100% training samples available to explore the effectiveness of each component with different amounts of data. Baseline model A extracts class names of anchor/target objects from $D$ using a language parser [6] and constructs the referential order according to the appearance of the names in $D$ directly. Also, we employ each of our Referring Blocks $R_{i}$ in model A without the FE module, i.e., directly using $F_{i}$ to calculate attention with text features of $O_{i:B}$ and $D$. Model B enhances the accuracy on 1% and 10% training data by conducting our two-stage object ordering with LLM. When further applying the order-aware pre-training in model C, significant improvements in all settings, especially for 1% available data, are observed. By conducting our pre-training strategy, our Vigor is able to learn foundational concepts of ordering and relations to locate the target objects progressively. Finally, our full model in the last row, incorporating FE modules each $R_{i}$ to enhance features of anchor/target objects with $F^{mask}_{i}$, achieves optimal results on both settings. This verifies the success of our proposed modules and warm-up strategy, especially when the available training data is very limited.
@@ 5 Conclusions
Acknowledgment.
This work is supported in part by the National Science and Technology Council via grant NSTC 112-2634-F-002-007 and NSTC 113-2640-E-002-003, and the Center of Data Intelligence: Technologies, Applications, and Systems, National Taiwan University (grant nos.113L900902, from the Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE) of Taiwan). We also thank the National Center for High-performance Computing (NCHC) for providing computational and storage resources.
Supplementary Material
@@ A.1 Pre-Training Sample Synthesis and Training Pipeline
This section provides the pseudocode of synthesizing order-aware samples for Vigor’s pre-training in Sec. 3.4 and the complete training pipeline in Sec. 3.5 in our main paper, respectively. Specifically, Algorithm 1 demonstrates the pipeline to synthesize an order-aware pre-training sample given object proposals $P$ and predicted object labels $L$. On the other hand, Algorithm 2 illustrates the complete pipeline to train Vigor with synthesized samples and natural-description samples, such as NR3D and ScanRefer.
Algorithm 1 Order-Aware Sample synthesis for Vigor Pre-training
@@ A.2 Implementation and Details and Experimental Settings
Algorithm 2 Vigor Training Pipeline
@@
6:     use 1 with $\{P,L,B\}$ as inputs to synthesize $\{D^{aug},O^{aug}_{1:B},p^{aug}_{1:B}\}$.
@@ A.2.1 NR3D
A.2.2 SR3D
We present SR3D’s results in App. B.2. SR3D contains 65844 training samples and 17726 testing samples. Each sample’s description is synthesized by simple spatial relations, such as farthest and beside with simple sentence structures like “the monitor that is farthest from the printer.” We conduct the object classification and warm-up process as in NR3D and continue on SR3D data pairs, with around 2.7k and 27k training steps for 1% and 100% data using batch size 24. SR3D 100% data training takes around 48 hours with one NVIDIA Tesla V100 GPU.
A.2.3 ScanRefer
For ScanRefer, following M3DRef-CLIP [56], we use PointGroup [22] to classify object proposals. We do not apply the warm-up due to noisy and imperfect object proposals and labels under ScanRefer’s setting. The batch size is 32, and the GPU is a single NVIDIA Tesla V100 GPU. Training on 100% data requires around 60 hours.
A.2.4 Baselines
For baselines [2, 17, 50, 21, 20, 5, 8, 57, 32, 56, 49] on NR3D and ScanRefer in Table 1,  2, and  3, we utilize their official public implementations with different amounts of available training samples and the full testing set to evaluate their low-resource performance. For the full-data ($100$%) scenario, we acquire their performance either on the official leaderboard of NR3D/ScanRefer or their published papers.
@@ B.1 More Results on NR3D
Figure 5: Quantitative results on NR3D We can see that when the amount of data is relatively many (above 30%), our Vigor is comparable to MVT+CoT3DRef [5]. However, as the amount of data reduces, our Vigor performs better over MVT-CoT3DRef.
Method Hard Easy View-Dep. View-Indep. Overall
Referit3D [2] 19.5 27.3 21.2 24.2 23.3
TransRefer3D [17] 21.6 29.9 22.9 27.0 25.7
SAT [50] 22.5 27.6 21.7 26.6 25.0
BUTD-DETR [21] 25.9 41.9 29.1 34.8 33.3
MVT [20] 22.9 30.3 25.4 27.1 26.5
MVT + CoT3DRef [5] 32.7 43.2 34.0 39.8 37.9
ViL3DRel + CoT3DRef [5] 32.4 44.7 33.4 40.9 38.4
Vigor (Ours) 39.1 53.3 45.3 46.4 46.0
In Sec. 4, we have shown comparisons between Vigor and several state-of-the-arts under data efficient scenarios (1% $\sim$ 10% of data). Here, we provide a comparison between Vigor and MVT+CoT3DRef [5] from 1% to 100% of NR3D data in Fig. 5, where Vigor outperforms CoT3DRef with a noticeable margin when the amount of data is limited and is comparable to CoT3DRef when the data amount is over 30%, showing that our Vigor is suitable under different settings.
B.2 More Results on SR3D
Table 5: Data Efficient Grounding accuracy (%) on SR3D. We show the results trained with 1% and 100% of training data.
@@ B.2 More Results on SR3D
1% 100%
BUTD-DETR [21] 36.5 67.0
MVT [20] 22.2 64.5
NS3D [18] 52.4 62.7
MVT + CoT3DRef [5] 26.9 73.2
Vigor (Ours) 51.3 67.1
Table 5 shows the quantitative comparisons on the SR3D dataset against BUTD-DETR, MVT, CoT3DRef and NS3D [18], with the settings of using $1\%$ and $100\%$ of training examples, respectively. From this table, we can see that although NS3D achieves the best performance when using only $1\%$ of training data due to its structured decomposition of the input description into nested logical expressions, its performance saturates when using more data for training. As for CoT3DRef, its design of applying rule-based matching of anchor/target objects as additional guidance for the model is effective on SR3D when using a large amount of data, where the relations are much simpler than NR3D or ScanRefer, but its performance on $1\%$ of data is suboptimal. On the contrary, our Vigor achieves comparable results in both settings, showing that our design is suitable for various amounts of training pairs.
@@ Appendix C Ablation Studies on Training Objectives
This section performs the ablation study on several training objectives used for Vigor mentioned in Sec. 3. In particular, we investigate the influence of $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ for low-resource (1% training data) and full-resource (100% training data) scenarios, as shown in Table 6. Note that we only conduct target object classification on the text feature $T$ when deactivating $\mathcal{L}_{text}$, and apply classification on both anchor and target objects when adopting $\mathcal{L}_{text}$, as mentioned in Sec. 3.3. Also, for all experiments in Table 6, the pre-training with synthetic data in Sec. 3.4 is applied for fair comparison. In summary, $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ all impose positive effects on models’ performances under the two settings, verifying Vigor’s design in Sec. 3.
Table 6: Ablation studies on training objectives. Note that for all ablation settings, the proposed order-aware pre-training in Sec. 3.4 is applied for fair comparison.
@@ Appendix D Model Performance Under Different Order Length
Figure 6: (a) Grounding accuracies with varying Object-Referring block numbers $B$ and (b) statistics of referential order length of NR3D. It can be seen that the model performance saturates at $B$=4, matching length statistics of NR3D.
Table 7: Grounding accuracy on NR3D subsets regarding different parsed referential order lengths. Note that referential order length $=1$ means only the target object is mentioned in the description. The improvement of Vigor over MVT [20] grows as the parsed order length increases.
@@
In this section, we analyze the effectiveness of padded order length $B$ of $O_{1:B}$ (and also the number of Referring Blocks) in Fig. 6, which illustrates the performances with different $B$ using 1% and 100% training data in Fig. 6(a). We also show the statistics of the original order length generated by our two-stage referential ordering in Fig. 6(b). We observe that although performances gradually increase with larger $B$ (i.e., considering more potential anchor objects appeared in $D$ achieves better accuracy), the performance gain saturates at $B=4$. This can be explained by looking at Fig. 6(b), which shows that only a very small amount of data exceeds an order length of $4$ in both training and testing pairs. This suggests that our selection of $B=4$ is reasonable, optimally balancing computational efficiency with prediction accuracy.
testing set) have 1 or 2 anchor objects mentioned other than the target object. Similarly, samples with an order length of 4 or 5 (209 samples in the testing set) have 3 or 4 anchor objects mentioned other than the target object. A longer referential order may generally denote a longer and more complicated description. As shown in Table 7, though only a 2.2% performance gain is obtained for target-only samples, Vigor is more advantageous when dealing with lengthy descriptions, with 5.3% performance gain achieved for descriptions with order length of 4 or 5.